OpenStudy (anonymous):

how can you solve 9^(3/2)?? without caluclator

OpenStudy (anonymous):

= (square root of 9) cubed

OpenStudy (anonymous):

= 3 ^3 = 27

OpenStudy (anonymous):

ok so for 4^(3/2) would be 8

OpenStudy (phi):

Yes. As you know \[(x^{a})^{b}= x^{ab}= (x^{b})^{a}\] so you can re-write \[4^{\frac{3}{2}}=(4^{\frac{1}{2}})^{3}\] and of course a power of 1/2 means take the square root, so we get 2^3= 8